Please Help!!!!!!!!!!!!

2(x + 1) = 4

Found 2 solutions by Alan3354, math1970:
Answer by Alan3354(69443)

2(x + 1) = 4
x + 1 = 2
x = 1
What else could it be?

Answer by math1970(1)

2(x + 1) = 4
2x + 2 = 4
2x + 2 - 2 = 4 - 2
2x = 2
2x/2 = 2/2
x = 1